# Pulse-PPG: An Open-Source Field-Trained PPG Foundation Model for Wearable Applications Across Lab and Field Settings

### v1

@@ 3.2.2. Exploiting the Noise of Real-World, mHealth PPG data:
(2) $d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{cand}})$ $\coloneqq\sum_{\textbf{x}_{\textrm{anc}}\in{\textbf{X}_{\textrm{anc}}}}\Bigg{(}\bigg{(}\sum_{\textbf{x}_{\textrm{cand}}\in{\textbf{X}_{\textrm{cand}}}}\underset{{\textbf{x}_{\textrm{cand}}\in{\textbf{X}_{\textrm{cand}}}}}{\textrm{softmax}}\Big{(}\textrm{sim}\big{(}f_{q}({\textbf{x}}_{\textrm{anc}}),f_{k}(\textbf{x}_{\textrm{cand}})\Big{)}f_{v}(\textbf{x}_{\textrm{cand}})\bigg{)}-\textbf{x}_{\textrm{anc}}\Bigg{)}^{2}$
@@
where $\smash{\textbf{X}\in\mathbb{R}^{T\times D}}$ and $\smash{\textbf{x}\in\mathbb{R}^{D}}$ with $T$ as the time length, and $sim()$ is a dot product. This function is implemented with a standard cross-attention block (Vaswani et al., 2017) with projection and normalizations layers been omitted for brevity.
@@
This work builds on (Xu et al., 2024a), who proposed a proof-of-concept motif-based distance for physiological signals in controlled, small-scale settings. However, their approach was limited to pre-training and evaluation on the same small labeled dataset, lacking the scalability and generalizability required for real-world foundation models. In contrast, our foundation model is pre-trained on a large, unlabeled field dataset and evaluated across diverse downstream tasks and datasets. Critically, we address noisy field conditions—a gap in prior work—by refining the methodology to handle real-world PPG variability, demonstrating robustness in dynamic, noisy environments.
3.2.3. Capturing Subtle Yet Meaningful Patterns:
Given our motif-based distance function, we can identify positive and negative pairs to draw together and push apart in the embedding space, respectively. Traditionally, contrastive approaches will utilize the Normalized-Temperature Cross-Entropy Loss (NT-Xent) function (Equation 4), in which the positive pair is pulled towards the anchor, and all negatives are pushed away, equally.
@@ 3.2.3. Capturing Subtle Yet Meaningful Patterns:
An alternative idea to address this is utilizing a metric learning loss function (Kim et al., 2019; Hoffmann et al., 2022; Kan et al., 2021), in which our self-supervised learning encoder learns to embed the PPG instances exactly as described by our motif-based distance function. However, although our learned motif-based distance function will be able to capture semantic information, it was trained without labels and is not expected to always, perfectly identify the correct semantic relationships.
A ”goldilocks” approach between these two ideas would be using a relative contrastive learning framework that models the relative positions of the distances. As such, we utilize the concept of relative dissimilarity via multiple negative sets (Xu et al., 2024b). For each positive pairing, a negative set is created by selecting samples whose distance from the anchor exceeds the distance of the positive pair. Each candidate in the dataset is iteratively used to form a positive pair, while the remaining candidates contribute to the negative set:
@@
4. Pre-training Procedure and Implementation for Pulse-PPG
@@ 4.1.1. Training the Learnable Motif-based Distance Function
In our initial label-free pre-training setting, we do not know if a random pair of PPG sequences share class labels, and thus we do not know if reconstruction error should be minimized to learn a motif-matching similarity function between them. Therefore, during training, we set ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$ to be the same value and apply a missingness mask on ${\textbf{X}}_{q}$. This mask has a contiguous segment of missingness that stretches 2 seconds. In this way, our distance function will learn to retrieve the regions from the key $\textbf{X}_{k}$ that match the missing 2 second motif from the query, $\bar{\textbf{X}}_{q}$, for reconstruction. Then, after training, we can utilize this distance as a static function for the RelCon loss to identify relative distances between ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$, where ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$ are different instances with different values.
@@ 4.1.2. Training the Pulse-PPG Foundation Model
  • Sampling within-subject, across-time, draws candidate PPG sequences from the same subject, but at a different time-window from the anchor. This allows us to model how an individual’s behavioral patterns throughout their day affects their PPG signal, for example via exercise.
  • Sampling between-subject draws candidate PPG sequences from other persons within a batch. This allows us to model how the similarities and differences in an individual’s specific physiology and routines affects their PPG signal.
@@ 4.2.1. Dataset Description:
Our pre-training dataset is composed of 822,247 unique 4-minute 50 Hz PPG segments, with 120 participants who wore a smartwatch a day for up to 100 days. This is from the Mobile Open Observation of Daily Stressors (MOODS) study dataset. The study was conducted using a WearOS app for smartwatches, a cross-platform smartphone app, and cloud services. Detailed descriptions of the System and Study Design of the MOODS Study are available in (Neupane et al., 2024).
4.2.2. Minimal Preprocessing:
To pre-train the physiological model, we utilize raw photoplethysmography (PPG) data collected from the participants. Out of 122 participants, the dataset recorded PPG data from 120 participants, which we use for pre-training. PPG data collected in the field is noisy, however, avoiding signal specific filtering can provide certain advantages for generalized, robust representations e.g., adapting to dataset diversity, preserving subtle meaningful variations, and transferability. Hence, we refrain from applying any particular filtering technique, but focus on addressing between person variability through a global person-specific z-normalization.
@@ 4.2.3. Subject-wise Train/Val/Test Splits:
To pre-train a model capable of handling the substantial within- and between-subject variability, we randomly shuffled participants and created a train/val/test split of 84/18/18. This split was consistently applied across all downstream tasks. We deliberately adopted a subject-based splitting strategy to pose a more challenging task for the model: making inferences on data from entirely unseen individuals, a common scenario in real-world applications.
@@ 4.2.4. PPG Sequence Inputs for Pre-Training Model
In our modeling approach, we use 4-minute-long windows as sequence inputs for the Pulse-PPG Foundation Model. This choice is grounded in research highlighting the cyclical nature of physiological responses to stress. Studies have shown that stress-induced patterns in heart rate and heart rate variability (HRV) follow cyclical behaviors, with durations typically ranging from 3.5 to 4.2 minutes (Lamichhane et al., 2017; Wilson et al., 2018; Bari et al., 2020). For example, heart rate increases during stress and remains elevated until the stressor ends, after which it returns to baseline, forming a cyclical pattern. Field studies, such as those involving stressful conversations or work-related stress, have further validated these patterns, with median cycle durations around 4 minutes (Bari et al., 2020).
By segmenting PPG signals into non-overlapping 4-minute windows, we aim to capture both micro-level perturbations (e.g., transient changes in pulse waveform morphology) and macro-level cyclical patterns (e.g., sustained heart rate elevation during stress). This approach allows our model to learn representations that are sensitive to the temporal dynamics of stress, which are critical for accurate stress detection. The 4-minute window size strikes a balance between capturing meaningful cyclical patterns and retaining fine-grained temporal details.
It is important to note that, although our initial pre-training inputs are 4 minutes long, our model incorporates a temporal pooling mechanism that enables it to generalize to inputs of varying lengths. This flexibility ensures that the model can be applied to diverse downstream tasks, where input durations may differ.
@@ 4.3.1. Encoder Details:
We utilize a 1D ResNet-26 network from the https://github.com/hsd1503/resnet1d code repository as our encoder backbone, with a total of 28.5m trainable parameters. There is instance normalization placed at the start of the ResNet, in order to account for PPG distribution shifts across users and across datasets. Our ResNet implementation utilizes an initial filter size of 128, a kernel size of 11, and a stride of 2, with 12 residual blocks that linearly increase the filter size every 4 blocks. There is a global average temporal pooling at the end of the network in order for our model to generalize across different time scales. As a result of this, our encoder embeds variable-length PPG signals into a single 1D 512-dimensional embedding vector.
@@ 4.3.2. Distance Function Details:
We utilize the code from the https://github.com/maxxu05/rebar code repository in order to implement the distance function. This is a lightweight model, used as a static function to identify relative distances for the RelCon loss function and only has 127k trainable parameters. The dilated convolution network is a series of a dilated convolution blocks, made up of a dilated convolution layer with residual connection skip, followed by a ReLU and instance norm. There are 5 blocks, with an initial dilation of 1 that doubles after each layer. The filter size is set to 64, the kernel size is set to 15, and stride is set to 10. The input convolution layer to the DilatedConvNet is set to be a partial convolution (Liu et al., 2018) in order to handle missingness in our pre-training and application procedures. The reconstruction error used as our distance function is a standard MSE loss.
4.3.3. Compute Details:
Our Pulse-PPG foundation model was trained for 5 days for 6 epochs, where each epoch is composed of 606,833 unique 4-minute PPG segments or about 5,424 unique participant days. The distance function was trained over 10 hours for 20 epochs. We utilized NVIDIA L40S GPUs for our experiments, and models were constructed with the Pytorch python package.
5. Results
The power of a foundation model lies in its ability to generalize across a broad range of tasks and datasets. This makes them especially valuable to the research community, providing pre-trained, general-purpose representations that can be easily adapted to new tasks, streamlining research efforts without the need for extensive task-specific data collection or manual feature engineering. Therefore, we design our experiments to validate this by evaluating our Pulse-PPG Foundation Model on 11 different downstream tasks, across 5 datasets, spanning typical wearable PPG tasks (i.e., stress classification) and typical hospital PPG tasks (i.e., systolic blood pressure regression). We evaluate this through a series of 3 experiments, 1) Pulse-PPG vs. A Prior PPG Foundation Model, 2) Assessing Field-to-Lab Generalizability, 3) Evaluation of Pulse-PPG Use-cases.
In the following sections we will describe our experimental design and results. In Section 5.1, we describe our downstream tasks and datasets. Then in Section 5.2, 5.3, and 5.4, we describe the design for each of our experiments and their results.

### v2

@@ 3.2.2. Exploiting the Noise of Real-World, mHealth PPG data
(2) $d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{cand}})$ $\coloneqq\sum_{\textbf{x}_{\textrm{anc}}\in_{s}{\textbf{X}_{\textrm{anc}}}}\Bigg{(}\bigg{(}\sum_{\textbf{x}_{\textrm{cand}}\in_{s}{\textbf{X}_{\textrm{cand}}}}\underset{{\textbf{x}_{\textrm{cand}}\in_{s}{\textbf{X}_{\textrm{cand}}}}}{\textrm{softmax}}\Big{(}\textrm{sim}\big{(}f_{q}({\textbf{x}}_{\textrm{anc}}),f_{k}(\textbf{x}_{\textrm{cand}})\big{)}\Big{)}f_{v}(\textbf{x}_{\textrm{cand}})\bigg{)}-\textbf{x}_{\textrm{anc}}\Bigg{)}^{2}$
@@
where $\smash{\textbf{X}\in\mathbb{R}^{T\times D}}$ and $\smash{\textbf{x}\in\mathbb{R}^{D}}$ with $T$ as the time length, $\in_{S}$ as a $\in$ but with a subsampling of stride $s$, and $sim()$ is a dot product. This function is implemented with a standard cross-attention block (Vaswani et al., 2017) with projection and normalization layers being omitted for brevity.
@@
This work builds on Xu et al. (2024a), which proposed a proof-of-concept motif-based distance for physiological signals in controlled, small-scale settings. However, their approach was limited to pre-training and evaluation on the same small labeled dataset, lacking the scalability and generalizability required for real-world foundation models. For example, the original work only evaluated on sequences with a maximum length of 3,840 time-points, much shorter than our 4-minute-long PPG input lengths of 12,000. This is significant because the backbone of the training function utilizes a cross-attention function, with a computational complexity of $O(T^{2})$, but we were able to address this with our $\in_{s}$ function with $s=10$. This allows for the reconstruction error to be calculated from subsample of the original sequence, effectively reducing computational complexity to $O((T/s)^{2})$, but without degrading the quality of motif similarity comparisons by allowing $f_{\{q/k/v\}}$ to capture the motifs found from the full sequence. Additionally, the prior work (Xu et al., 2024a) utilized a simple, untuned dilated convolution encoder to act as a control to fairly compare self-supervised learning methods. In our work, we utilize a much more complex ResNet encoder backbone to develop a powerful foundation model that can capture and handle the noise and diversity in real-world PPG data. To this end, it necessitated an extensive tuning of the methodology (see Section 4.3.1) to work in our setting, but due to this, we have been able to demonstrate robustness in the dynamic field environment across diverse downstream tasks and datasets.
3.2.3. Capturing Subtle Yet Meaningful Patterns
Given our motif-based distance function, we can identify positive and negative pairs to draw together and push apart in the embedding space, respectively. Traditionally, contrastive approaches will utilize the Normalized-Temperature Cross-Entropy Loss (NT-Xent) function (Equation 4), in which the positive pair is pulled towards the anchor, and all negatives are pushed away equally.
@@ 3.2.3. Capturing Subtle Yet Meaningful Patterns
An alternative idea to address this is utilizing a metric learning loss function (Kim et al., 2019; Hoffmann et al., 2022; Kan et al., 2021), in which the encoder learns to embed the PPG instances exactly as described by our motif-based distance function. However, although our learned motif-based distance function will be able to capture semantic information, it was trained without labels and is not expected to always perfectly identify the correct semantic relationships.
A ”goldilocks” approach between these two ideas would be using a relative contrastive learning framework that models the relative positions of the distances. As such, we utilize the concept of relative dissimilarity via multiple negative sets (Xu et al., 2024c). For each positive pairing, a negative set is created by selecting samples whose distance from the anchor exceeds the distance of the positive pair. Each candidate in the dataset is iteratively used to form a positive pair, while the remaining candidates contribute to the negative set:
@@
This work builds on Xu et al. (2024c), which proposed relative contrastive learning to learn a foundation model for accelerometry sensor data. However, accelerometry sensor data varies significantly from PPG sensor data. For example, accelerometry has a well-established set of invariances (i.e., 3D rotation) (Tang et al., 2020), due to how inertial its measurements are fundamentally governed by physics—the magnitude and temporal patterns of acceleration reflect biomechanical forces generated by movement, regardless of sensor orientation or placement. PPG signals, by contrast, exhibit no such geometric invariances—their morphology is highly sensitive to sensor contact, motion artifacts, and physiological variability. Rather than presupposing geometric symmetries, for Pulse-PPG, we allow the motif-based distance function to implicitly learn noise-robust representations directly from the raw PPG waveform’s temporal structure. Additionally, their work, as well as other accelerometry machine learning work, demonstrates that short time windows can capture semantically meaningful motion information, and as such, they only encode accelerometry segments spanning 2.56 seconds, or 256 time points. However, for Pulse-PPG, we seek to model 4-minute (12,000 time point) PPG segments, to align with prior stress models (Lamichhane et al., 2017; Wilson et al., 2018; Bari et al., 2020). Therefore, this also results in an extensive tuning of the methodology (see Section 4.3.1) to work in our setting. For example, we reduce the same-subject candidates from 20 to 1 and constrain them to be drawn from the same hour as the anchor. Due to this tuning, we have been able to develop Pulse-PPG, a powerful foundation model that generalizes across many different PPG settings and tasks.
4. Pre-training Procedure for Pulse-PPG
@@ 4.1.1. Training the Learnable Motif-based Distance Function
In our initial label-free pre-training setting, we do not know if a random pair of PPG sequences shares class labels, and thus, we do not know if reconstruction error should be minimized to learn a motif-matching similarity function between them. Therefore, during training, we set ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$ to be the same value and apply a missingness mask on ${\textbf{X}}_{q}$. This mask has a contiguous segment of missingness that stretches 2 seconds. In this way, our distance function will learn to retrieve the regions from the key $\textbf{X}_{k}$ that match the missing 2-second motif from the query, $\bar{\textbf{X}}_{q}$, for reconstruction. Then, after training, we can utilize this distance as a static function for the RelCon loss to identify relative distances between ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$, where ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$ are different instances with different values.
@@ 4.1.2. Training the Pulse-PPG Foundation Model
  • Sampling within-subject, across-time, draws candidate PPG sequences from the same subject, but at a different time window from the anchor. This allows us to model how an individual’s behavioral patterns throughout their day affect their PPG signal, for example, via exercise.
  • Sampling between-subject draws candidate PPG sequences from other persons within a batch. This allows us to model similarities and differences in physiology among different individuals.
@@ 4.2.1. Dataset Description:
For our pre-training, we obtained access to the Mobile Open Observation of Daily Stressors (MOODS) study dataset (Neupane et al., 2024) that is composed of 822,247 unique 4-minute 50 Hz PPG segments, with 122 participants who wore a smartwatch every day for up to 100 days. Participants included 39 men, 77 women, and 6 non-binary individuals, with an average age of 38 years ($SD=13$). Seventy-one were working professionals across fields such as academia, information technology, administration, public safety, health, law, sales, and construction; fifty were students pursuing higher education, and one participant did not disclose their occupation. Participants wore a study-provided Fossil Sport (Version 4) smartwatch, and approximately 55K hours of raw PPG data, sampled at 100 HZ, were collected during the study. The study was conducted using a WearOS app for smartwatches, a cross-platform smartphone app, and cloud services. Detailed descriptions of the system and study design of the MOODS study are available in (Neupane et al., 2024).
@@ 4.2.2. Subject-wise Train/Val/Test Splits:
To pre-train a model capable of handling substantial within- and between-subject variability, we randomly shuffled participants and created a train/val/test split of 84/18/18. During pre-training, the test set is never used, however, we establish this split so that the downstream tasks that utilize this same dataset (i.e., Wearable Field Stress and Activity) have a held-out test set never exposed during pre-training. We deliberately adopted a subject-based splitting strategy to pose a more challenging task for the model: making inferences on data from entirely unseen individuals, a common scenario in real-world applications.
4.2.3. Preserving Noisiness:
To pre-train the physiological model, we utilize raw photoplethysmography (PPG) data collected from the participants. Out of 122 participants, the dataset recorded PPG data from 120 participants, which we use for pre-training. PPG data collected in the field is noisy; however, avoiding signal-specific filtering can provide certain advantages for generalized, robust representations, e.g., adapting to dataset diversity, preserving subtle meaningful variations, and better transferability. Hence, we refrain from applying any particular filtering technique, but focus on addressing between-person variability through a global person-specific $z$-normalization. Please see Appendix A.3 for our ablation study, which shows that normalization consistently enhances performance.
@@ 4.2.4. PPG Sequence Inputs for Pre-Training Model
In our modeling approach, we use 4-minute long windows as sequence inputs for the Pulse-PPG Foundation Model. This choice is grounded in research highlighting the cyclical nature of physiological responses to stress. Studies have shown that stress-induced patterns in heart rate and heart rate variability (HRV) follow cyclical behaviors, with durations typically ranging from 3.5 to 4.2 minutes (Lamichhane et al., 2017; Wilson et al., 2018; Bari et al., 2020). For example, heart rate increases during stress and remains elevated until the stressor ends, after which it returns to baseline, forming a cyclical pattern. Field studies involving stressful conversations or work-related stress, have further validated these patterns, with median cycle durations around 4 minutes (Bari et al., 2020).
By segmenting PPG signals into non-overlapping 4-minute windows, we aim to capture both micro-level perturbations (e.g., transient changes in pulse waveform morphology) and macro-level cyclical patterns (e.g., sustained heart rate elevation during stress). This approach allows our model to learn representations that are sensitive to the temporal dynamics of stress, which are critical for accurate stress detection. The 4-minute window size strikes a balance between capturing meaningful cyclical patterns and retaining fine-grained temporal details. Please refer to Appendix A.3 for the ablation study of evaluating the effects of varying input window sizes. Overall, the 4-minute window yields the best performance, consistently outperforming the 1- and 2-minute models across most metrics and ranking within the top two.
It is important to note that, although our initial pre-training inputs are 4 minutes long, our model incorporates a temporal pooling mechanism that enables it to generalize to inputs of varying lengths. This flexibility allows our model to be utilized on our diverse downstream tasks, in which the input windows are variable in length.
@@ 4.3. Pulse-PPG Model Implementation Details
4.3.1. Model Adaptation and Tuning
As described in Sections 3.2.2 and 3.2.3, we adapt the recent motif-similarity-based reconstruction and relative contrastive loss-based self-supervised modeling approach (Xu et al., 2024b, c) so that it can capture the intricacies of 4-minute segments of PPG data.
We began with a parameter search for the distance function model, which was responsible for motif-based reconstruction of masked PPG segments. In this phase, we experimented with different kernel sizes (5, 11, 15, 21, 60, 120), embedding dimensions (32, 64, 128), double receptive field values (to control how many layers and how much temporal context the network could attend to), and learning rates (0.001, 0.0001). Smaller kernel sizes and shallower receptive fields failed to capture sufficient temporal context for motif alignment, while overly large fields led to over-smoothing and poor reconstruction on noisy segments. The selected configuration provided the best trade-off between learning localized motifs and capturing broader temporal dependencies in field data.
Once the distance model was trained, we used it to train a 1D ResNet self-supervised encoder. In our parameter exploration for this ResNet encoder, we evaluated different kernel sizes (9, 11, 15), initial base filter sizes (32, 64, 128), stride sizes (1, 2, 5, 10, 15, 20), number of ResNet blocks (4, 8, 12, 16), number of within-user same-hour candidates (1, 5, 10, 15), final pooling strategies (mean, max), and learning rates (0.001, 0.0001). We found that while deeper architectures marginally improved performance on training subsets, they generalized poorly on held-out field data, likely due to overfitting on participant-specific noise. Pooling strategies also played a significant role: max pooling consistently produced more discriminative embeddings for downstream tasks, especially when separating subtle variations in physiological activity. Once we identified the best combination that gave us reasonable performance on the 10-day dataset, we initiated the same process for the full 100-day dataset, using a more informed combination of parameters rather than a random search. In this stage, we repeated the parameter search process for both learning stages and found the best hyperparameter combinations to be:
  • Distance learning: kernel size = 15, embedding dim = 64, double receptive field = 5, epochs = 20, learning rate = 0.001, batch size = 16, and stride = 10.
  • Contrastive learning: base filters = 128, kernel size = 11, stride = 2, groups = 1, nblocks = 12, final pool = max, epochs = 20, learning rate = 0.0001, and batch size = 64.
@@ 4.3.2. Encoder Details:
We utilize a 1D ResNet-26 network from the https://github.com/hsd1503/resnet1d code repository as our encoder backbone, with a total of 28.5m trainable parameters. There is instance normalization placed at the start of the ResNet to account for PPG distribution shifts across users and datasets. Our ResNet implementation utilizes an initial filter size of 128, a kernel size of 11, and a stride of 2, with 12 residual blocks that linearly increase the filter size every 4 blocks. There is a global average temporal pooling at the end of the network for our model to generalize across different time scales. As a result of this, our encoder embeds variable-length PPG signals into a single 1D 512-dimensional embedding vector.
@@ 4.3.3. Distance Function Details:
We utilize the code from the https://github.com/maxxu05/rebar code repository to implement the distance function. This is a lightweight model, used as a static function to identify relative distances for the RelCon loss function, and only has 127k trainable parameters. The dilated convolution network is a series of dilated convolution blocks, made up of a dilated convolution layer with a residual connection skip, followed by a ReLU and instance norm. There are 5 blocks, with an initial dilation of 1 that doubles after each layer. The filter size is set to 64, the kernel size is set to 15, and the stride is set to 10. The input convolution layer to the DilatedConvNet is set to be a partial convolution (Liu et al., 2018) to handle missingness in our pre-training and application procedures. The reconstruction error used as our distance function is a standard MSE loss.
4.3.4. Training Details:
Our Pulse-PPG foundation model was trained for 5 days for 6 epochs on an NVIDIA L40S GPU. Each epoch is composed of 606,833 unique 4-minute PPG segments or about 5,424 unique participant days, and due to this large data size, each epoch took 20 hours to finish. We chose this early stopping point due to convergence of the pre-training loss, (2) diminishing returns on downstream task performance, and (3) computational efficiency—extending further would yield negligible benefits while significantly increasing the carbon footprint. This can be seen in the training and validation loss curves visualized in Figure 5. Both training and validation loss are calculated for each batch, but aggregated across the entire epoch for reporting. The training loss is backpropagated and weights are updated after each step for each batch (accelerating convergence and lowering computational cost), aligning with deep learning standards (Amari, 1993). An Adam optimizer was used with a learning rate set to 0.0001, betas set to (0.9, 0.999), no weight decay, and a batch size of 64.
The distance function was trained over 10 hours for 20 epochs with an Adam optimizer with a learning rate of 0.001, betas set to (0.9, 0.999), no weight decay, and a batch size of 16. Models were constructed with the PyTorch Python package, and full reproducibility details can be found in our final released codebase.
5. Evaluation Set-up
The power of a foundation model lies in its ability to generalize across a broad range of tasks and datasets. This makes them especially valuable to the research community, providing pre-trained, general-purpose representations that can be easily adapted to new tasks, streamlining research efforts without the need for extensive task-specific data collection or manual feature engineering. Therefore, we design our experiments to validate this by evaluating our Pulse-PPG Foundation Model on 11 different downstream tasks, across 5 datasets, spanning typical wearable PPG tasks (i.e., stress classification) and typical hospital PPG tasks (i.e., systolic blood pressure regression). In the following sections, we will describe our evaluation set-up and experimental results. We describe our downstream datasets and tasks in Section 5.1, evaluation methodologies in Section 5.2, evaluation metrics in Section 5.3 and experimental results in Section 6.